Circular No. 7855

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 F1 (UTSUNOMIYA) Preliminary parabolic orbital elements from MPEC 2002-F39:

T = 2002 Apr. 23.460 TT  ω = 124.^{o}093
Ω = 288.167 }2000.0
q = 0.45854 AU  i = 78.800
2002 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Mar. 17 21^{h}37.^{m}90 + 3^{o}55.^{′}9 1.659 0.979 32.^{o}5 33.^{o}1 9.5
19 21 45.17 + 5 31.8 1.614 0.943 32.6 34.7 9.3
21 21 52.92 + 7 12.8 1.570 0.907 32.6 36.3 9.1
23 22 01.21 + 8 59.0 1.526 0.871 32.6 38.0 8.8
25 22 10.11 +10 50.3 1.484 0.835 32.4 39.8 8.6
27 22 19.71 +12 46.6 1.444 0.800 32.1 41.5 8.3
29 22 30.10 +14 47.6 1.406 0.764 31.7 43.3 8.1
31 22 41.40 +16 52.6 1.369 0.729 31.1 45.1 7.8
Apr. 2 22 53.72 +19 00.7 1.336 0.695 30.5 46.9 7.5
4 23 07.19 +21 10.2 1.305 0.661 29.7 48.6 7.3
6 23 21.94 +23 19.4 1.278 0.629 28.9 50.2 7.0

V838 MONOCEROTIS T. R. Geballe, Gemini Observatory; A. Evans, B. Smalley, and V.

H. Tyne, Keele University; and S. P. S. Eyres, University of Central Lancashire, report J, H, and K spectroscopy of V838 Mon from Mar. 9, obtained as on IAUC 7796: “The previously reported (IAUC 7796, 7829) deep CO absorption bands longward of 2.29 μm are now in emission and consist of a strong and almost smooth elevated ‘plateau’ on which are superimposed narrow absorptions at the individual ^{12}CO band heads, observable at least up to v = 8–6. Weaker dips at the wavelengths of ^{13}CO band heads also are present. The water bands previously reported (IAUC 7796) are no longer apparent, and the strong absorption feature near 1.09 μm is much weaker. There is now stronger Paβ emission at 1.282 μm than previously, and Brγ (2.166 μm) is more weakly in absorption; we see Brackett lines from γ to 4–24, but with no significant changes in line strengths compared with Jan. observations. The P-Cyg profiles previously reported on metal lines are now much less pronounced. The overall flux level has increased by a factor of ∼ 15 since our last spectrum (IAUC 7796), but the 1–2.5-μm continuum is still well approximated by a ∼ 4000-K Planck function.”

2002 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green